How many grams of chlorine gas must react to give 3.72 g of BiCl3?
2Bi(s)+3Cl2(g)→2BiCl3(s)
Express your answer with the appropriate units.
1
Expert's answer
2020-12-01T09:31:40-0500

2 BiCl3 = 3 Cl2

2(315.34g) BiCl3 = 3(70.9 g) Cl2

1g BiCl3 = 3 ×70.9 /(2×315.34)g Cl2

3.72g = (3 ×70.9 × 3.72) /(2 ×315.34g) Cl2

= 1.25g